# Answer, Refuse, or Guess? Investigating Risk-Aware Decision Making in Language Models

### v1

@@ 4.3 Implementation Details
Figure 8: Proportions of expected-value reasoning.
4.4 Main Results
We report scores for six LMs across three datasets (MedQA, MMLU, and GPQA) under four representative risk levels in Table 2. The reported scores are calculated by averaging the performance across three prompt variations. Here are the key findings:
4.4.1 Prompt Chaining Consistently Boost Performance in High-Risk Settings
In the high-risk columns $(1,-8)$ and $(1,-4)$, every LM shows a notable performance gain over other three methods when using the prompt chaining strategy. Nearly all LMs reach their best scores, with the only one exception of gemini-1.5-pro-002 on MMLU. This indicates that decomposing the task into separate inference steps for each skill is especially beneficial in high-risk scenarios.
4.4.2 Stepwise Prompts Do NOT Offer Consistent Improvements
In high-risk settings, although the stepwise prompt (Figure 6, left) explicitly outlines the three skills, it seldom outperforms the simpler risk-informing prompt (Figure 3). By contrast, prompt chaining consistently yields better results. This comparison highlights that merely decomposing skills does not always guarantee a performance boost; the method to implement skill decomposition is also crucial. We discuss why prompt chaining is more effective than stepwise prompts in section 5.1.
4.4.3 Risk-Informing Methods Degrade Performance in Low-Risk Settings
Under low-risk settings $(4,-1)$ and $(8,-1)$, risk-informing methods—including risk-informed prompts, stepwise prompts, and prompt chaining—actually reduce the average performance compared to the baseline (no risk informed) approach. We discuss the implications of this finding in section 5.2.
5 Discussion
5.1 LMs’ Ability to Reason with Expected Values Is Elicited by Prompt Chaining
As suggested in Section 3, our core hypothesis is that LMs’ expected-value reasoning ability remains hidden when the compositional task also requires other skills. To test whether skill decomposition elicits this ability, we examined how often LMs explicitly calculate expected values to answer or refuse. Figure 7 illustrates examples of using (left) versus not using (right) expected-value reasoning.
To quantify the results, we provided gpt-4o-mini-2024-07-18 with a 6-shot prompt containing three positive (using expected-value reasoning) and three negative (lacking such reasoning) examples, and instruct it to classify LMs’ output. As shown in Figure 8, prompt chaining dramatically increases the proportions of explicit expected-value calculations, whereas risk-informing or stepwise prompts do not. Notably, simply scaling the models does not consistently enhance proportions of expected-value reasoning for risk-informing or stepwise prompts. We further discuss why prompt chaining is more effective than stepwise prompts in Appendix D.
Figure 9: A scatter plot of $\Delta$performance versus refusal.
5.2 Why Does Prompt Chaining Perform Worse in Low-Risk Settings?
Prompt chaining underperforms the “no risk informed” baseline in low-risk settings. We assume it is because LMs still refuse some questions, thereby losing the positive expected reward from random guessing. To test this, we plot each experiment run as a dot on Figure 9. The three dots on the lower right, for instance, represent running three prompt variations with gpt-4o-mini on GPQA. The figure shows that higher refusal rates correlate with greater degradation. This finding gives a straightforward solution: In low-risk scenarios with positive expected rewards, do not give LMs the option to refuse–let them answer every question instead.
@@ 5.2 Why Does Prompt Chaining Perform Worse in Low-Risk Settings?
5.3 Do Reasoning Models Eliminate the Need for Skill Decomposition?
@@
Recent progress in reasoning LMs–such as o1, o1-mini, o3-mini, and DeepSeek-R1–has substantially improved performance on complex reasoning tasks Guo et al. (2025). This naturally raises the question of whether these models can already tackle risk-aware decision making without skill decomposition. Due to the high cost of reasoning LMs and our budget constraints, we run o3-mini-2025-01-31 and DeepSeek-R1 on GPQA in high-risk settings as representative experiments, and show the results in Table 3. Results indicate that reasoning LMs still require skill decomposition through prompt chaining to handle the task more effectively.
6 Related Work
Enabling Language Models to Refuse
LMs tend to hallucinate when they lack sufficient knowledge to fulfill instructions Ji et al. (2023), so increasing efforts have been made to enable them to refuse. Zhang et al. (2024) identify queries that exceed LMs’ parametric knowledge and construct refusal-aware instruction-tuning data to fine-tune them. Feng et al. (2024) propose a collaborative approach where multiple LMs probe each other to detect knowledge gaps, thereby improving refusal accuracy. Xu et al. (2024) use reinforcement learning with a reward model to incentivize refusal for out-of-knowledge questions. Our work extends this line of research by examining not just whether an LM should refuse, but also when it should do so in the context of varying risk levels.
Biases of LMs in Decision Making
Prior work discusses decision-making biases of LMs from various perspectives. Horton (2023) replicates classic experiments in behavioral economics Kahneman et al. (1986); Samuelson and Zeckhauser (1988); Charness and Rabin (2002) on GPT-3 Brown et al. (2020), showing that it often reproduces human-like biases. Ross et al. (2024) further study LMs’ behavioral tendencies in economic games via utility theory. Jia et al. (2024) propose an evaluation framework to estimate an LM’s risk preference, probability weighting, and loss aversion. However, these studies primarily rely on toy datasets or game-based scenarios, focusing on measuring biases rather than identifying their causes or offering solutions. In contrast, we identify LMs’ issues from a compositional generalization perspective, and propose skill-decomposition methods to address the identified shortcomings.
Compositional Generalization
Lake and Baroni (2018) introduce the SCAN dataset to study compositional generalization (CG) in neural networks, revealing that they may lack compositional abilities. Since then, this topic has received increasing attention. For instance, Li et al. (2021) study CG in encoder-decoder Transformers for machine translation. Press et al. (2023) investigate LMs’ performance gap between compositional and individual questions on multi-hop question answering. Hosseini et al. (2024) examine LMs’ instance-level CG in arithmetic questions ($Q$), where each instance is composed in a way that a variable of $Q_{2}$ depends on the solution to $Q_{1}$. This work extends the study of CG to a new task of risk-aware decision making, revealing new compositional gaps in current LMs.
7 Conclusion and Future Work
In this work, we formalize the task of risk-aware decision making, and reveal current LMs’ suboptimal behaviors across different risk levels. We draw two main conclusions from the experiments: (1) In high-risk settings, skill decomposition via prompt chaining substantially improves performance for both regular and reasoning LMs. (2) In low-risk settings, we should simply omit the refusal option and allow LMs to attempt every question, given the positive expected reward of guessing.
By framing risk numerically, we can systematically evaluate LMs’ performance and identify their irrational decision-making behaviors. However, real-world risk often appears through more subtle natural-language cues, like the patients’ health conditions in Figure 1. How LMs’ biases manifest in these less structured scenarios remains an open question, which we leave for future work.
Limitations
Risk Representation
In our evaluation framework, we presents risk levels to LMs as numeric values, which enables us to conduct controlled experiments to study LMs’ behaviors systematically. However, risk might be described in natural language in practical scenarios. For example, in medical practice, health conditions are typically documented in textual descriptions. For patients with cormobidities, the consequences of making the wrong decision are severe, which corresponds to the high-risk setting; for patients with healthy underlying conditions, the consequences would be more benign, which corresponds to the low-risk setting. In these scenarios, either humans must formulate risk factors into numeric scores, or LMs themselves must learn to integrate textual descriptions of risk into the decision-making process. How to quantify LMs’ performance and biases in these scenarios remains an open question.
Improvements on Individual Skills
In our skill-decomposition solution, we decompose the task of risk-aware decision making into three skills: downstream task, confidence estimation, and expected-value reasoning. Therefore, one obvious approach to boost the final performance is to improve these individual skills. For example, there are a lot of techniques to improve LMs’ confidence estimation from existing calibration studies Guo et al. (2017); Lin et al. (2022); Tian et al. (2023); Xiong et al. (2024). Applying these techniques to the confidence estimation step has the potential to enhance the overall performance in risk-aware decision making, as evidenced by the discussion in Appendix E, where we show that better confidence estimation indeed contributes to higher scores. As a result, although we do not elaborate how to improve performance on these individual skills in detail, we consider this as a promising direction.
Integrating Existing Refusal Mechanisms
Our work complements existing refusal strategies by extending a new research direction on how LMs should adapt their decisions according to varying risk levels. In principle, existing methods Feng et al. (2024); Xu et al. (2024); Zhang et al. (2024) could be integrated in our framework to further refine refusal performance. However, we view these two lines of research (i.e., refusal mechanisms per se vs. risk-aware decision making) as orthogonal, so we do not pursue a detailed integration of their methods in this paper. Future research might explore a unified system that leverages both existing refusal techniques and risk-aware decision making to develop safer, more reliable LMs.
Appendix A Prompt Templates
@@
A.1 Paraphrased Prompts for Pure Gambling
@@
In the pure gambling experiment (see section 3.2), we paraphrase the initial pure gambling prompt in Figure 5 into 100 variations in order to assess LMs’ refusal rates. Here we list three of the variations in Figure 12, 13, and 14 for better understanding of how the paraphrased prompts look like.
A.2 Prompt Variations for Main Experiments

### v2

@@ 3.1 Suboptimal Behaviors in Language Models’ Decision-Making Policies
3.2 Isolating the Ability of Handling Expected Values
If we think of the skills required for LMs to maximize expected reward on a given task, there are at least three individual types of skills:
Downstream task: The skills required to solve the given task. In our work, they’re the knowledge and reasoning skills of various domains (e.g., biology, physics, and chemistry on GPQA).
Confidence estimation: The LM’s ability to estimate how likely its answer is correct. This ability is often referred to as calibration in the literature (Guo et al., 2017; Lin et al., 2022; Kadavath et al., 2022; Tian et al., 2023; Xiong et al., 2024).
Reasoning with expected values: Based on the estimated confidence, calculate the expected reward or penalty to guide the answer-or-refuse decision. For example, if the LM is 95% confident in answering the question correctly and $(r_{\mathrm{cor}},r_{\mathrm{inc}})=(1,-8)$, the reasoning process might look like Figure 4 (left). We refer to this skill as expected-value reasoning.
To investigate why LMs fail to apply expected-value reasoning when answering questions, our first step is to remove the need for the other two skills and directly evaluate LMs’ ability to apply expected-value reasoning. We design a “pure gambling” experiment to isolate this skill and observe two things: (1) whether LMs apply expected-value reasoning more frequently than they do when solving multi-choice questions, and (2) whether their refusal proportions become more aligned with the optimal decision policies.
Figure 5 displays the prompt for the pure gambling experiment. Since we cannot assess refusal rates using a single prompt instance, we paraphrased the prompt into 100 variations with equivalent semantic meaning (e.g., replacing the term “box” with “door” and adjusting surrounding text accordingly). Examples of the paraphrased prompts are shown in Appendix B.2. Because the pure gambling task still provides four choices, our previously defined high-risk (negative expected reward) and low-risk (positive expected reward) settings still apply.
Like what we did in subsection 3.1, we sampled 100 instances and manually inspected LMs’ outputs to see whether they apply expected-value reasoning to guide their decision. We found that they apply expected-value reasoning in 95 instances, which is a striking difference from only 4 instances when LMs are answering multi-choice questions. As for the refusal proportions, Figure 3(b) reports the refusal proportions for each LM under this pure gambling scenario. Compared to results in Figure 3(a), we observe more optimal decision policies in both high-risk and low-risk scenarios. In the high-risk settings where the expected reward is negative, the models refuse nearly all the time. For example, when $(s_{\mathrm{cor}},s_{\mathrm{inc}})=(0,-1)$, the 6-42% refusal rate in Figure 3(a) improves to 93-99% in Figure 3(b). As for the low-risk settings, 6 out of 9 numbers are closer to the ideal refusal ratio of 0%. Especially, when $(s_{\mathrm{cor}},s_{\mathrm{inc}})=(1,0)$, all tested LMs exhibit the optimal policy of zero refusal.
Given these findings, a clear distinction emerges: LMs demonstrate a strong capability for applying expected-value reasoning when presented with a “pure” decision problem, as in the pure gambling scenario. However, this type of reasoning is not reliably activated when it is not the only required skill, but rather a meta-level component of task that requires additional knowledge and reasoning, like scientific question answering in GPQA. The suboptimal decision policies observed in subsection 3.1, therefore, do not seem to be the lack of individual skills, but an inability to autonomously compose them within a single inference process. This insight motivates our next section.
Pure gambling prompt There are four identical boxes. Only one box contains a prize, and the other three boxes are empty. Choices: A. Open box A B. Open box B C. Open box C D. Open box D The scoring criteria is as follows: If you open the box with the prize, you get $r_{\mathrm{cor}}$} point(s). If you open an empty box, you get $r_{\mathrm{inc}}$} point(s). If you choose not to open any box, you get $r_{\mathrm{ref}}$} point(s). Your goal is to score as high as possible based on the above scoring criteria, and you can either choose to: (1) Open a box by outputting a single letter (A, B, C, or D) (2) Choose not to open any box by outputting a single letter N
@@ 3.2 Isolating the Ability of Handling Expected Values
Figure 5: The pure gambling prompt to evaluate the skill of expected-value reasoning.
4 Improving Risk-Aware Decision Making
Based on our finding that LMs fail to consistently apply expected-value reasoning to guide their risk-aware decision making when they need to compose different skills, we now investigate strategies to improve their performance. Our investigation is centered around the following two questions:
Q1: Compared to not informing LMs of the risk, does explicitly informing an LM of the risk really improve its ability to maximize expected reward?
Q2: Can we further improve performance by decomposing the risk-aware decision-making task into explicit steps, guiding the model to apply each required skill sequentially?
To answer these questions, we evaluate the following four distinct prompting strategies, which represent a progression from the standard baseline to more human-guided interventions.
Figure 6: The stepwise prompt (left) versus prompt chaining (right).
No-Risk Prompt (Baseline). This is a standard multiple-choice QA prompt (Figure 7) that omits information about the risk structure or the option to refuse, representing typical LM usage.
Risk-Informing Prompt. (Figure 2) It explicitly provides the risk structure $(r_{\mathrm{cor}},r_{\mathrm{inc}},r_{\mathrm{ref}})$ and the goal of maximizing the total score. Compared this method to the baseline directly addresses Q1.
Stepwise Prompt. To address Q2, this method decomposes the task into sequential steps within a single prompt. As illustrated in Figure 6 (left), it explicitly guides the model to first solve the problem, then estimate its confidence, and finally use that confidence to make a risk-aware decision, all within a inference pass.
Prompt Chaining. This method (Figure 6, right) also decomposes the task but uses multiple inference steps. The output from each step is programmatically extracted and passed as input to the next step, ensuring the LM to apply each skill separately.
A critical aspect of LM evaluation is prompt sensitivity. Following the recommendations of Mizrahi et al. (2024), we conduct a multi-prompt evaluation to enhance the robustness of our findings. Specifically, for each of our four methods, we create three prompt variations (See Appendix B.3), all semantically identical but phrased differently. We compare the performance of these four methods using the normalized reward metric $R$ (averaged by three prompts) in low-risk and high-risk settings.
4.1 Low-Risk Settings
We argue that the answers to Q1 and Q2 are rather obvious in low-risk settings. Based on our discussion in subsection 2.2, we know that random guessing yields a positive expected reward ($r_{\mathrm{guess}}>r_{\mathrm{ref}}=0$), so the optimal policy is to always answer. The no-risk prompt (baseline), lacking a refusal option, naturally enforces this optimal behavior. In contrast, the other three methods introduce the possibility of refusal, which is theoretically a suboptimal decision. As expected, we empirically confirm that the no-risk prompt (baseline) performs best in low-risk settings (see Appendix C for full empirical results), so we focus our main analysis on the high-risk scenario.
Table 1: Performance ($R$) of four methods under $(r_{\mathrm{cor}},r_{\mathrm{inc}})=(1,-8)$
Method/Model C-Haiku GPT-mini Gem-Flash C-Sonnet GPT-4o Gem-Pro Average
Dataset MMLU
No-risk prompt -0.723 -0.668 -0.624 -0.004 -0.115 -0.335 -0.412
Risk-informing -0.704 -0.651 -0.572 0.121 -0.076 -0.283 -0.361
Stepwise prompt -0.762 -0.683 -0.648 0.197 -0.169 -0.273 -0.389
Prompt chaining -0.305 -0.016 -0.175 0.307 0.327 -0.066 0.012
Dataset MedQA
No-risk prompt -1.062 -0.818 -1.370 0.050 0.147 -0.661 -0.619
Risk-informing -1.026 -0.789 -1.325 0.064 0.159 -0.520 -0.573
Stepwise prompt -1.034 -0.835 -1.414 0.083 0.166 -0.565 -0.600
Prompt chaining -0.835 -0.083 -0.537 0.324 0.475 -0.316 -0.162
Dataset GPQA
No-risk prompt -4.669 -4.391 -3.699 -2.885 -3.336 -2.957 -3.661
Risk-informing -4.507 -4.273 -3.618 -2.503 -3.210 -2.717 -3.471
Stepwise prompt -4.502 -4.334 -3.815 -1.756 -3.294 -2.901 -3.434
Prompt chaining -3.129 -0.969 -1.178 -0.413 -0.054 -2.127 -1.312
Table 2: Performance ($R$) of four methods under $(r_{\mathrm{cor}},r_{\mathrm{inc}})=(1,-4)$
Method/Model C-Haiku GPT-mini Gem-Flash C-Sonnet GPT-4o Gem-Pro Average
Dataset MMLU
No-risk prompt 0.043 0.073 0.094 0.442 0.380 0.257 0.215
Risk-informing 0.060 0.084 0.120 0.480 0.367 0.290 0.233
Stepwise prompt 0.029 0.046 0.080 0.471 0.317 0.311 0.209
Prompt chaining 0.165 0.295 0.199 0.504 0.444 0.309 0.319
Dataset MedQA
No-risk prompt -0.146 -0.010 -0.318 0.472 0.526 0.076 0.100
Risk-informing -0.116 0.014 -0.293 0.455 0.537 0.144 0.123
Stepwise prompt -0.166 -0.173 -0.329 0.466 0.531 0.118 0.075
Prompt chaining -0.073 0.267 -0.060 0.490 0.565 0.159 0.224
Dataset GPQA
No-risk prompt -2.166 -1.998 -1.619 -1.158 -1.416 -1.201 -1.593
Risk-informing -2.081 -1.938 -1.559 -1.057 -1.412 -1.069 -1.519
Stepwise prompt -2.099 -1.972 -1.693 -1.479 -1.552 -1.222 -1.670
Prompt chaining -1.621 -0.435 -0.607 -0.892 -1.115 -0.964 -0.939
4.2 High-Risk Settings
A more interesting and non-trivial case is the investigation of high-risk settings, where $r_{\mathrm{guess}}<r_{\mathrm{ref}}=0$ so LMs must answer selectively by weighing potential reward and penalty. Since the optimal policy is no longer straightforward, we empirically evaluate the performance of the four methods with six different LMs from three model families, with implementation details provided in Appendix E. Results are shown in Table 1 and Table 2. The answer to Q1 is rather coherent: the risk-informing method consistently outperforms the no-risk prompt across six models, three datasets, and two risk structures. That is, informing LMs of the risk indeed help them achieve the goal of maximizing expected reward. On the other hand, the answer to Q2 is more nuanced: prompt chaining performs the best in almost all combinations (except for gemini-1.5-pro-002 on MMLU), while stepwise prompt does not always provide improvement over the risk-informing prompt. This shows that while skill decomposition indeed improves LMs’ ability to maximize expected reward, it comes with a subtlety: one should instruct LMs to apply one skill at a time with separate inference steps. The reason behind this may be related to the curse of instructions: where Harada et al. (2024); Jaroslawicz et al. (2025) found that LMs cannot reliably follow multiple instructions at once. In our case, We also conduct experiments to show that the performance of each individual skill is indeed worse in stepwise prompt compared to prompt chaining (see Appendix F).
@@ 4.2 High-Risk Settings
Do Reasoning Models Eliminate the Need for Skill Decomposition? Recent progress in reasoning LMs has substantially improved performance on complex reasoning tasks (Guo et al., 2025). This naturally raises the question of whether these models can already tackle risk‐aware decision making without skill decomposition. Due to the high cost of reasoning LMs and our budget constraints, we run o3-mini-2025-01-31 and DeepSeek-R1 on GPQA in high-risk settings as representative experiments, and show the results in Table 3. Results indicate that reasoning LMs still require skill decomposition through prompt chaining to handle the task more effectively.
5 Discussion and Conclusions
Advantages and Disadvantages of Skill Decomposition. There is a particular advantage of skill decomposition through prompt chaining: one can apply existing techniques to improve performance of the individual skills. We hereby list several examples for each skill. For solving the downstream task, one may spend more inference-time compute to achieve better answer quality with methods like self-consistency (Wang et al., 2022), best-of-N (Cobbe et al., 2021; Snell et al., 2024), or budget forcing (Muennighoff et al., 2025) (for reasoning LMs specifically). For confidence estimation, there are a lot of related works on how to improve LMs’ calibration such as temperature scaling (Guo et al., 2017; Tian et al., 2023), probing the last-layer hidden state (Zhang et al., 2025), or training LMs to verbalize confidence better (ConfTuner) (Li et al., 2025). We also conduct a small study to show that improving calibration also boosts the final performance metric $R$ in Appendix G. For the final step of expected-value reasoning, an LM may call external tools (Schick et al., 2023) (e.g., calculators) if it is not good at arithmetic or to save inference cost.
However, the primary disadvantage is increased overhead. Prompt chaining requires sequential inference calls, raising both latency and computational costs. Practitioners must therefore weigh the trade-off between better decision-making performance and higher costs.
Main Takeaways. Our work introduces a systematic framework for evaluating risk-aware decision-making and uncovers a critical flaw in current LMs: they consistently deviate from optimal policies. Specifically, they tend to over-answer in high-risk scenarios, incurring avoidable penalties, and over-defer in low-risk settings, giving up positive expected rewards from random guessing. The main cause behind such failure is not lack of knowledge of expected-value reasoning, but from an inability to apply this knowledge reliably when LMs need to compose multiple skills for reward maximization. Our findings also offer straightforward guidance. For low-risk applications where expected reward is positive, one may consider to enforce the LM to always answer. Conversely, for high-risk applications, reliability can be enhanced by using skill decomposition through prompt chaining to guide the model through a structured process of solving the downstream task, estimating confidence, and making a final, risk-weighted decision.
From a broader perspective, our findings highlight a critical gap on the path to truly autonomous agents. While current LMs possess a remarkable range of individual skills, they struggle to spontaneously compose them reliably. The risk-aware decision-making setting investigated here is just one instance of this compositional failure. It remains an open question how many other types of tasks exist where LMs will falter, potentially requiring an endless series of human-designed scaffolds like skill decomposition. This suggests that the road to building agents that can autonomously reason and act reliably in novel scenarios is still long.
Ethics Statement
The development of risk-aware decision-making capabilities in language models represents a significant step toward more reliable AI systems, but it also raises important safety and ethical considerations. While our research aims to improve AI safety by enabling models to defer appropriately under uncertainty, there are substantial risks if these capabilities are misinterpreted or misapplied. Organizations may use improved risk calibration as justification for autonomously deploying AI systems in high-stakes domains such as medical diagnosis, financial advisory, or legal decision-making, where human expertise and oversight remain essential regardless of technical improvements. More broadly, enhanced risk-aware decision-making might create a false sense of security that encourages premature automation in consequential applications, undermining human agency and accountability in critical decisions. Such misuse threatens not only individual safety in specific applications but also public trust in AI systems and the responsible development of autonomous technologies. As language models continue to advance in their decision-making capabilities, it is crucial to emphasize that benchmark improvements do not substitute for domain-specific validation, regulatory compliance, and maintained human oversight. In conducting this research, we focus exclusively on established public benchmarks, provide transparent evaluation frameworks, and emphasize that our findings represent progress toward safer AI rather than validation for autonomous deployment in consequential real-world applications.
Reproducibility Statement
All experiments using closed source LLMs were conducted on using the latest released python package by each vendors ( OpenAI, Google Gemini ). For closed source LLMs we opted using Together.AI hosted LLMs service for the results. All models were used using the best recommended configurations from their respective papers or official releases. Approximately \$2,000 USD was spent on API access for closed and open weight models. Code will be released on publicly accessible platforms (GitHub) upon submissions.
Appendix A Use of LLMs in Writing
We used ChatGPT and Gemini to improve the writing of paper, mainly on detecting grammar errors or mis-spellings.
Appendix B Prompt Templates
@@
B.1 No risk informed prompt
Figure 7 shows the standard zero-shot CoT prompt commonly used for answering multiple-choice questions.
No risk informed prompt Answer the following question based on the provided choices: Question: {question} Choices: {choices_text} Provide your reasoning process first, then provide your final answer in the following format: ANSWER: \$letter
Figure 7: The regular prompt which does not inform the LMs of the risk.
B.2 Paraphrased Prompts for Pure Gambling
@@
In the pure gambling experiment (see section 3.2), we paraphrase the initial pure gambling prompt in Figure 5 into 100 variations in order to assess LMs’ refusal rates. Here we list three of the variations in Figure 8, 9, and 10 for better understanding of how the paraphrased prompts look like.
Figure 8: The first example of a paraphrased pure gambling prompt.
Figure 9: The second example of a paraphrased pure gambling prompt.
Figure 10: The third example of a paraphrased pure gambling prompt.
B.3 Prompt Variations for Main Experiments